OpenStudy (anonymous):

a bobsled slides down an ice track starting (at zero initial speed) from the top of a(n) 134 m high hill the acceleration is 9.8 m/s2. neglect friction and air resistance and determine the bobsleds speed at the bottom of the hill. answers in units of m/s.

OpenStudy (jusaquikie):

v^2=v0^2 +2a d v^2=0+19.6(134) v^2=2626.4 v=51.24